OpenStudy (anonymous):

If i have 19 out of 39 correct what is my percentage?

OpenStudy (bahrom7893):

39-100 19-x x= 19*100/39

OpenStudy (bahrom7893):

that's the right percentage

OpenStudy (anonymous):

#corect/#total move the decimal to the right 2 spaces, 48.7%

OpenStudy (anonymous):

(19/39)*100 =48.71

OpenStudy (anonymous):

Any of these 3 ways will get you the same result.

OpenStudy (anonymous):

using a calculator compute \[19\div 39\] then using your eyeballs move the decimal two places right

OpenStudy (anonymous):

no one in real life multiplies by 100. just move the decimal

OpenStudy (anonymous):

^ thats what i was thinking hah

OpenStudy (bahrom7893):

LOL I do!